Show complete solution. Use equation function.

1. Find the general solution of ada = 9x²dx- ydy.

2. Find the particular solution of y' = ycosx/1+2y² given y (x) = 1; y (0) = 1.

3. Find the general solution to the equation (cos y + 2) y' = 2x.


1
Expert's answer
2021-06-01T12:20:54-0400

(1)."ada=9x^2dx-ydy"

Integrating both sides:


"\int ada= \int9x^2dx -\int ydy"


"\dfrac{a^2}{2}=9 \times\dfrac{x^3}{3} -\dfrac{y^2}{2} +C"


"\dfrac{a^2}{2}= 3{x^3} -\dfrac{y^2}{2} +C"


(2). "\dfrac{dy}{dx} =\dfrac{y cosx}{1+2y^2}"


On variable separation we get:


"(\dfrac{1}{y} +2y)dy=cos xdx"


Integrating both sides:


"\int(\dfrac{1}{y} +2y)dy= \int cos xdx"


"log y +y^2=sinx +C"


Now it is given that:

y=1 when x=0, so put in above equation:


"log(1)+1=sin(o) +c"


"C=1"

So the solution is:


"logy+y^2=sinx +1 \Rightarrow logy+y^2-sinx -1 =0"


(3)."(cosy+2) \dfrac{dy}{dx}=2x"


"cos ydy+2dy=2xdx"

"siny+2y =x^2+c"

"siny+2y -x^2=c"